Ben throws a rock at an angle 45° from horizontal on level ground. He throws the rock with a velocity of 30m/s. Create a acc vs. t graph, a v vs. t graph and a displacement vs. time graph for the x and y directions.
1
Expert's answer
2020-04-08T10:53:15-0400

As the rock is thrown at an angle 45 then there will be two component of velocity

one being the "v_x=30\cos45^o" and "v_y=30\sin45^o"

as acceleration is only in y direction then "a_x=0" and "a_y=-10m\/s^2"

hence graph of acceleration will be a straight line for for y and zero for x direction


let's see for velocity

v=u+at

for x : "v_x=u_x=15\sqrt2(a_x=0)"


for y :"v_y=u_y+a_yt=15\sqrt2-10t"





let's see for displacement

s=ut+0.5at2

for x: "S=15\sqrt2t+0 =15\sqrt2t"



for y: "S=15\sqrt2t-5t^2"